OpenStudy (anonymous):

Want a medal? I need to create 3 linear equations for three objects. A baseball, a basketball, and a football helmet. I don't know where to begin. Then I need to create a Literal Inequality using roller blades? I know I need to create variables for the objects so I guess I'll use the first letter but I don't know how to creatively make linear equations for each of these objects..Please help<3 This isn't multiple choice or anything I just need to create literal equations with all variables in the formula in the correct positions. Final solution is needs to be correct also

OpenStudy (triciaal):

just lost the post computer issues

OpenStudy (triciaal):

try again the short version my idea is to think of a sports store variables baseball.b basketball, n football, f rollaeblades, r assign cost for each example 10, 4, 2 budget constraint $300 to spend and need to get 25 items b + n + f = 25 10b + 4n + 2f </= 300

OpenStudy (anonymous):

Thank you but it needs to be 3 individual linear equations @triciaal

OpenStudy (triciaal):

wait the store only has helmets and rollerblades how many in store example h + r = 55 r = 55 - h or h = 55 - r r = 55 - 25 = 30 how this gives an idea

OpenStudy (greencat):

u cud make quadratics for the heights of each.

OpenStudy (greencat):

oops linear /bonk

OpenStudy (triciaal):

I need to create 3 linear equations for three objects.

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (triciaal):

another idea using the cost for specific combination of items eg 2b + 3f+ 5n = 200 3b + 2f+ n = 150 b + n + f = 30

OpenStudy (greencat):

speed of baseball vs how hard it was thrown strength of helmet vs amount of material used to make it

OpenStudy (anonymous):

@ganeshie8